OpenStudy (lynda619):

Can someone help me with these two questions? 4.8x^2=8 x= 2.56x^2=20 x=

OpenStudy (welshfella):

4.8 x^2 = 8 divide both sides by 4.8:- x^2 = 8 / 4.8 = 5/3 s = +/- sqrt (5/3)

OpenStudy (welshfella):

x = sqrt 5/3 , - sqrt 5/3

OpenStudy (welshfella):

you can do the second one by the same method